# VHE gamma-ray intranight variability from BL Lacertae during the extreme flaring state of 2022

### 25 Mar 2026 v2

@@ 2.4 Ultraviolet data: Swift-UVOT
The analysis of UVOT data was carried out using the uvotimsum and uvotsource tasks within the HEAsoft 6.32 package, along with the latest (v20240201) Swift/UVOT Calibration Database (CALDB) files. UVOT performed photometric observations on BL Lac using the optical vv, bb, uu and UV w1, m2, w2 filters (Poole et al., 2008). Through uvotsource, we estimated the observed fluxes in the UVOT filters using a circular source region with radius of 5 centred on BL Lac and a circular background region with a 10 radius in a nearby region devoid of sources. The observed fluxes were corrected for Galactic extinction effects using (Schlafly & Finkbeiner, 2011) using the standard procedure described in Fitzpatrick (1999). Moreover, the contributions of the host galaxy in the UVOT photometric filters vv, bb, uu, w1, m2, w2, given the adopted aperture, were estimated as 2.89, 1.30, 0.36, 0.026, 0.020, and 0.017 mJy, respectively (Raiteri et al., 2013). Both host galaxy contributions and Galactic extinction effects were accounted for to reconstruct the intrinsic BL Lac spectral points in the optical-UV band. The results of the analysis of the Swift-UVOT data from the 2022 high state are reported in Table 7.
@@ 2.5 Optical data
Optical observations in the Johnson-Cousins bands were taken by the Sierra Nevada Observatory (SNO) as part of the TOP-MAPCAT photo-polarimetric blazar monitoring programme (Agudo et al., 2012). The data were reduced with the IOP4 software (Escudero Pedrosa et al., 2024a, b). Observations in the band were also performed as part of the Tuorla blazar monitoring444https://tuorlablazar.utu.fi/ with the 80 cm Joan Oró Telescope at Montsec Astronomical Observatory (Spain), reduced as detailed in Nilsson et al. (2018). Data in the bands were taken by the IAC80 telescope at the Teide Observatory. Finally, observations with the robotic telescope network of Las Cumbres Observatory (Brown et al., 2013) were also coordinated in the optical bands.
Additionally, data from the public all-sky survey programmes and databases All-Sky Automated Survey for Supernovae (ASAS-SN, Shappee et al., 2014; Kochanek et al., 2017) in the Sloan band, and Zwicky Transient Facility (ZTF, Bellm et al., 2019) in the Sloan bands, were also included. These Sloan magnitudes were converted into the Johnson-Cousins photometric system using the transformations from Lupton (2005)555https://classic.sdss.org/dr4/algorithms/sdssUBVRITransform.php#Lupton2005.
The contribution of the host galaxy emission to the total observed optical emission contained in the employed aperture has been subtracted (see Raiteri et al., 2009). Moreover, all data have been corrected from the reddening effect introduced by the Galactic extinction. For this, we adopt the extinction values , , and reported by Schlafly & Finkbeiner (2011). The optical flux density light curves are presented in Fig. 1.
@@ 2.6 Radio-millimetric data
Observations in the 37-GHz radio band were performed with the 13.7-m diameter Metsähovi radio telescope as part of its long-term AGN monitoring666https://www.aalto.fi/en/metsahovi-radio-observatory/active-galaxies (Teraesranta et al., 1998). Additionally, we compiled five observations in the 225-GHz band performed with the Submillimeter Array (SMA, Ho et al., 2004) at Mauna Kea, Hawaii. SMA is an 8-element 6-meter diameter dish radio interferometer with two orthogonally polarised receivers, allowing us to measure the total and the polarised flux of astrophysical objects. These receivers are inherently linearly polarised, however they can be converted to circular thanks to quarter-wave plates installed on the polarimeter (Marrone & Rao, 2008). These observations were carried out by the SMA Monitoring of AGNs with Polarisation (SMAPOL) programme (Myserlis et al., 2025), dedicated to the regular monitoring of -ray loud blazars, including BL Lac, on a bi-weekly cadence. The data were calibrated with the MIR777https://lweb.cfa.harvard.edu/~cqi/mircook.html package following standard procedures.
@@ 3 VHE intranight variability
In addition to the day-to-week variability, we have evaluated the VHE -ray intranight variability, a feature less often observed in blazars (other examples are e.g. Aharonian et al., 2007; MAGIC Collaboration et al., 2020). Its detection presents significant constrains to the emitting region size and the interpretation of the emission, potentially favouring leptonic models as hadrons usually require longer times for being accelerated and have longer cooling times by a factor . We evaluate the nights with the highest VHE -ray flux, October 20 and November 13. The 5-minute binned light curves are shown in Fig. 2.
3.1 October 20th flare
We have evaluated the fastest variability timescales observed during the night of October 20. For this night we consider the integrated flux above an energy of 100 GeV, as all the data were taken under dark-sky conditions and zenith distances 20-40∘. As shown in the left panel of Fig. 2, the rise and decay of the flare is well sampled by the LST-1 data, varying from 1.0 C.U. to 4.4 C.U. in hours. This well defined rise and decay profile of the flare during this night enabled us to model the light curve as a rising and decaying exponential function, commonly used for modelling light curves (Norris et al., 1996) including those from blazar flares (see e.g. Aharonian et al., 2007),
where represents the flux of the flare at a time , is the baseline flux that we have fixed at cm-2 s-1, obtained from the stable bins at the beginning and end of the flare, and and are the timescales of the flare rise and decay variations by a factor of . We have modelled the 5-minute binned light curve of October 20 as a sum of three possible substructures during the development of the flare with peak values found at min, min and min, considering the first flux point as min. We note that we have also considered the simpler cases of one or two components, comparing the goodness of the fit in each case. We measure values for the triple-, double- and single-exponential profiles of , and , which yields a significance of the triple-exponential model over the double- and single-function of and , respectively, using Wilks theorem (Wilks, 1938). We also checked that the light curve modelling does not improve adding a fourth component, with a preference over the three-component one of 0.2. Therefore, we adopted the model based on the sum of three exponential profiles defined by Eq. (1).
The fitted function is shown in the left panel of Fig. 2, and the calculated rise and decay timescales of exponential profile are reported in Table 1. These timescales can be used to estimate the doubling flux variability timescale during the rise and decay, representative of the fastest significant variability timescales, as
With this approach we are able to resolve variability timescales of 18 min with a confidence level of ¿, and faster hints of between 2.2 min and 4.6 min, however with a significance 2.
@@ 3.1 October 20th flare
Alternatively, we can also estimate the fastest variability timescales through the doubling time criterion considering flux variations between consecutive temporal bins, following the procedure from Zhang et al. (1999) as
where and are the fluxes at times and . We consider only those timescales obtained through this method for which the significance of the variability is 2. The results derived with this approach are reported in the last column of Table 1. We observe significant doubling flux timescales of min.
Applying causality arguments, it is expected that very fast variability originates in very compact jet regions. Therefore, the fastest variability timescales observed can be used to constrain the size of the emitting region responsible for such variations as
where is the radius of the emitting region, represents the fastest variability timescale observed, is the speed of light, is the Doppler factor of the region and corresponds to the redshift of the source (see Dondi & Ghisellini, 1995; Tavecchio et al., 2010). Here we adopt a conservative approach, using the shortest timescale with a significance ¿ before trials, this is, min. With this timescale, and considering a range of Doppler factor values , the size of the emitting region responsible for the fast variability must be cm to cm. We note that if the 2.2- to 6.7-minute timescales are confirmed, this would imply a region a factor 1.5 to 4 more compact. Nevertheless, in addition to the low significance, when considering the typical systematic errors for LST-1 flux measurements (see Abe et al., 2023), these timescales get hindered within uncertainties.
3.2 November 13th flare
The same analysis was performed for the night of November 13, when we also observed intranight variability in the VHE -ray emission of BL Lac. In this case, and owing to the higher energy threshold introduced by the moon conditions, we consider the flux above 200 GeV. During this night, the emission decayed from 2.8 C.U. to 1.0 C.U. in 1.5 hours. The corresponding 5-minute binned light curve is shown in the right panel of Fig. 2. In this case, LST-1 covered only the decay of the VHE -ray flare, that we modelled with a decaying exponential function as
that follows the same description as Eq. (1). A value of cm-2 s-1 was used for , where the VHE flux seems to stabilise after the flare. The results for this night are reported as well in Table 1. With the estimated timescale min, we have again constrained the size of the emitting region with Eq. (4), for typical Doppler factors . This leads to a size constraint ranging from cm to cm. We note that in this case, the doubling flux criterion does not find significant minute-scale variability.
@@ 4 Characterisation of the HE-VHE -ray spectrum
In order to evaluate possible -ray spectral variability patterns, we have studied the evolution of the -ray spectrum during the flare. For this, we have applied a Bayesian Block (BB) analysis on the VHE -ray light curve using the algorithm developed by Wagner et al. (2022)888https://github.com/swagner-astro/lightcurves, identifying 10 BBs of comparable -ray activity with a false positive rate of 1% (see Fig. 1). Then, making use of the LAT and LST-1 data, we characterise the high-energy SED peak for the different identified periods from 100 MeV up to a few TeV. For this, we perform a joint fit of the data from both instruments with Gammapy, that allows to combine multi-instrument data at an event level, correctly accounting for instrument response functions (IRFs), backgrounds and other effects such as EBL absorption (Nievas Rosillo et al., 2025).
@@
For the Fermi-LAT data of each BB, we filter out all sources in the model that have a TS¡10. The parameters of all sources contained in a circular region of 10∘ radius centred at the position of BL Lac and with a TS¿30 are left free, while for those sources within 5∘ from BL Lac with TS¡30, only the spectral normalisation is left as a free parameter. For the remaining sources, all parameters are fixed to the catalogue values. Then, a pre-fit of the Fermi-LAT data is performed, allowing to remove weak sources still included in the model for which the number of expected counts is ¡5. Finally, we combine the LAT and LST-1 datasets of each BB and calculate the complete -ray spectrum and SED peak, incorporating into the analysis the response of each instrument.
@@
We have tested several spectral shapes of the HE-VHE emission, including log-parabola, power-law with exponential cut-off and log-parabola with exponential cut-off spectral models. We find that the first function provides the most accurate fit to our data. Therefore, we use a model consisting on a log-parabola function plus a term accounting for the EBL absorption at redshift with the model from Saldana-Lopez et al. (2021),
@@
To perform a direct comparison between the different spectra, we adopted an GeV for all periods. The best-fit spectral parameters are reported in Table 2. In Fig. 3 we show an example of the joint SED for the eighth BB, that is, the night of November 13. The rest of the SEDs are available in Figs. 6 and 7. Alternatively, owing to the very small variability in the spectral curvature, we tested a log-parabola function fixing , taken from the average value over the entire period. The results were consistent within uncertainties with those reported in Table 2.
@@
The complete characterisation of the high-energy SED peak performed in Sect 4 allows us to search for spectral variability patterns during the evolution of the flare. We observe a relatively stable value of the spectral index at 1 GeV around , with almost all BBs consistent within errors at a 2 level. Nevertheless, spectral index variations are particularly noticeable for the nights of October 20 with , and November 13 with , meaning that during these flares, particles are being accelerated to much higher energies. This is also visible from the high-energy peak shift, reported in Table 2. The IC peak shift becomes evident for the four brightest BBs, this is, the nights of October 20 and 27, and November 13 and 25, being the only ones with GeV. The average value of is found below 2 GeV, and during these nights we observe shifts up to values of 5.5 GeV for October 20, 6.5 GeV for November 13, and 2.3 GeV for October 27 and November 25. The Spearman correlation coefficient between and , and and show a high level of correlation ( and , respectively), indicating a possible harder-when-brighter behaviour. However, this correlation coefficient is dominated by the two brightest nights and does not account for uncertainties.
@@ 5 Broadband SEDs
We have modelled the evolution of the broadband emission of BL Lac during the 2022 high state. We have considered 10 broadband SEDs, corresponding to each of the BBs identified above. Fig. 4 shows as an example the model for the eighth BB. All fitted SED models are presented in Figs. 8 and 9. We note that, due to the lack of strictly simultaneous Swift data (0.5 days with respect to the LST-1 observations) to BBs 3, 4, 6, 8 and 9, for these periods we have considered a window of 1.5 days in order to have complete multi-wavelength coverage of the SEDs.
@@
We have considered a leptonic-only origin for the broadband emission. We use version v0.4.0 of the python package agnpy (Nigro et al., 2022, 2023). We find that SSC models based on a single emitting region are not able to accurately describe the broadband emission of BL Lac during this period. One-zone models based on EC scattering rise as an alternative to explain the VHE -ray emission, and have actually been invoked in the past for BL Lac due to the presence of a small and weak broad line region (BLR), using its external low-energy photons as seeds for the -ray production (MAGIC Collaboration et al., 2019). However, this solution was also found to be insufficient in this case to reproduce the broadband emission and variability during the 2022 high state, especially given the tight constrains on the emitting region size set by the intranight variability. Therefore, we consider a SSC model based on two emitting regions (Tavecchio et al., 2011), which has been used frequently to model the multi-wavelength emission of this source (see e.g. MAGIC Collaboration et al., 2019; Sahakyan & Giommi, 2022). We denote the two regions as ”core” and ”blob”, with sizes . We consider the regions to be co-spatial and interacting. This is motivated by the high correlation of the broadband emission from the optical band up to rays (Appendix D.2). A scheme of this layout is illustrated in Fig. 9b of MAGIC Collaboration et al. (2019). Each region is defined by a magnetic field , a Lorentz factor , a Dopper factor and an electron population that we have characterised through a broken power-law distribution with respect to the electron Lorentz factor ,
@@
where is the normalisation of the distribution, and are the minimum and maximum values of the Lorentz factor, and and are the spectral indices below and above the break of the distribution , respectively. We constrain the size of the blob, responsible for the intranight VHE -ray variability, with the estimates presented in Sect. 3. Additionally, we have incorporated to the SED model an EC contribution from external photon fields, namely the BLR, the accretion disc and the dusty torus. This is because, as mentioned above, these components have been used in the past to reproduce the large width of the IC SED peak (see e.g. Khatoon et al., 2024). Following Ghisellini & Tavecchio (2009); MAGIC Collaboration et al. (2019), we adopt erg s-1 and cm for the BLR, which leads to erg s-1 and cm using the typical relations between BLR, disc and torus (see Kaspi et al., 2007; Cleary et al., 2007), assuming in this case a covering factor for torus of , and a BLR reprocessing factor (based on the Hα line) . For the dusty torus, we assumed a typical temperature of 1000 K. We adopted a black hole mass of (Ghisellini et al., 2010). Finally, we considered the core located at a distance cm from the central engine and the blob at cm. We also take into account the possible absorption at GeV-TeV energies and the EBL absorption at VHE. We also evaluated the absorption due to the external radiation fields, which was found to be negligible.
@@
Multi-zone models of blazar SEDs are characterised by a high degree of degeneracy, owing to the relatively large number of parameters. This degeneracy generally prevents the use of standard fit routines based on the minimisation of a cost function to effectively constrain the parameter space. In such cases, it is necessary to adopt a physically motivated scenario as guiding framework. In this work, we have investigated a scenario in which the evolution of the flare is driven by magnetic reconnection, which is strongly supported by the minute-scale variability detected at VHE with LST-1. In order to fit the SED of each BB, for each block we have used the parameters of the previous one as initial parameters, iterating until reaching a good representation of the SED. To reduce the number of free parameters, we have kept the magnetic fields , the emitting region sizes , the Lorentz factor and the Doppler factor fixed (with values in agreement with the limit values set by the variability analysis). Additionally, we minimise the changes of physical parameters in the core region between adjacent BBs to reflect the slow evolving nature of the core, unless the observed time interval between the centre of two adjacent BBs is longer than two days, the light crossing timescale of the core region in our reference frame. The blob, in contrast, is considered as the newly formed plasmoid from an unresolved magnetic reconnection layer in the core, whose parameters can vary considerably depending on the reconnection layer. Therefore, we have based our SED evolution in changes of the electron distribution of each region, that is, , , , , and . The parameter change was done in a way that is consistent with the flare’s evolution timescales. The resulting models for each BB are shown in Figs. 8 and 9, and the parameters describing each of them are summarised in Table 3. The consistency of these models with a magnetic reconnection scenario is discussed in Sect. 6.
@@
The leptonic origin of the emission considered here is supported by recent results obtained by the IXPE (Imaging X-ray Polarimetry Explorer) satellite, based on the analysis of the broadband polarised emission of BL Lac (see Middei et al., 2023; Peirson et al., 2023). Under a leptonic framework where the X-ray spectrum is dominated by the high-energy IC scattering component, the X-ray polarisation degree is expected to be substantially lower than that of the optical band, which originates from electron synchrotron radiation (Liodakis et al., 2019; Peirson & Romani, 2019; Zhang et al., 2024). Middei et al. (2023) measured an upper limit of % for two IXPE observations of BL Lac during May and July 2022, consistent with this scenario given the optical polarisation degree.
@@
Moreover, under strong flares where the soft X-ray emission has significant contribution from synchrotron radiation as observed from the much flatter X-ray spectrum — which is also the present case, see for instance BB3 with respect to e.g. BB8 when the source was at its brightest — one expects a significantly higher polarisation degree of the soft X-ray photons than that at hard X-ray energies, and even higher than the optical polarisation degree (Liodakis et al., 2019; Peirson & Romani, 2019; Zhang et al., 2024). A third IXPE observation of BL Lac, performed two days after the last LST-1 observation (November 27 to 29) resulted in a measured X-ray polarisation degree of over the entire X-ray band. After splitting the analysis in two bins accounting for the soft and hard X-ray regimes, almost all the polarised emission was found to be in the former, with a polarisation degree of . The synchrotron emission was estimated to comprise 44% of the total flux at the soft X-ray band (Peirson et al., 2023). After accounting only for the synchrotron emission and assuming that the 56% of the flux remaining comes from unpolarised SSC, the X-ray polarisation fraction was calculated to be 49%, much higher than that in optical frequencies (13%), and well within the expectations for a SSC scenario (Peirson & Romani, 2019). Therefore, the leptonic SED modelling performed here is consistent with the expected nature of the broadband emission based on both the spectral and the multi-wavelength polarisation analyses.

### 07 Apr 2026 v3

@@ 2.4 Ultraviolet data: Swift-UVOT
We analysed the UVOT data using the uvotimsum and uvotsource tasks within the HEAsoft 6.32 package, along with the latest (v20240201) Swift/UVOT calibration database (CALDB) files. The UVOT obtained photometric observations on BL Lac through the optical vv, bb, uu and UV w1, m2, w2 filters (Poole et al., 2008). We estimated the observed fluxes in the UVOT filters using uvotsource, adopting a circular source region with a radius of 5 centred on BL Lac and a circular background region with a 10 radius in a nearby source-free region. We corrected the observed fluxes for Galactic extinction effects using (Schlafly and Finkbeiner, 2011), following the standard procedure described in Fitzpatrick (1999). We estimated the host galaxy contributions in the UVOT photometric filters vv, bb, uu, w1, m2, and w2, given the adopted aperture, to be 2.89, 1.30, 0.36, 0.026, 0.020, and 0.017 mJy, respectively (Raiteri et al., 2013). We accounted for both the host galaxy contribution and Galactic extinction effects to reconstruct the intrinsic BL Lac spectral points in the optical-UV band. Table 7 lists the Swift-UVOT data from the 2022 high state.
@@ 2.5 Optical data
We obtained optical observations in the Johnson-Cousins bands at Sierra Nevada Observatory (SNO) as part of the TOP-MAPCAT photo-polarimetric blazar monitoring programme (Agudo et al., 2012). We reduced the data using the IOP4 software (Escudero Pedrosa et al., 2024a, b). We also obtained observations in the band as part of the Tuorla blazar monitoring programme444https://tuorlablazar.utu.fi/ using the 80 cm Joan Oró Telescope at Montsec Astronomical Observatory (Spain), and reduced the data as detailed in Nilsson et al. (2018). Additional data were obtained in the bands with the IAC80 telescope at the Teide Observatory. We coordinated observations in the optical bands with the Las Cumbres Observatory robotic telescope network (Brown et al., 2013).
We also included data from the public all-sky survey programmes and databases: the All-Sky Automated Survey for Supernovae (ASAS-SN; Shappee et al., 2014; Kochanek et al., 2017) in the Sloan band, and the Zwicky Transient Facility (ZTF; Bellm et al., 2019) in the Sloan bands. We converted these Sloan magnitudes into the Johnson-Cousins photometric system using the transformations from Lupton (2005)555https://classic.sdss.org/dr4/algorithms/sdssUBVRITransform.php#Lupton2005.
We subtracted host galaxy contributions to the observed optical emission within our aperture (see Raiteri et al., 2009), and we corrected all data for Galactic extinction reddening. For this, we adopted the extinction values , , , and reported by Schlafly and Finkbeiner (2011). Fig. 1 presents the optical flux density light curves.
@@ 2.6 Radio-millimetric data
The 13.7-m diameter Metsähovi radio telescope obtained 37-GHz radio band observations as part of its long-term AGN monitoring programme 666https://www.aalto.fi/en/metsahovi-radio-observatory/active-galaxies (Teraesranta et al., 1998). Additionally, we compiled five observations in the 225-GHz band, obtained with the Submillimeter Array (SMA; Ho et al., 2004) at Mauna Kea, Hawaii. The SMA is an eight-element 6-metre-diameter dish radio interferometer with two orthogonally polarised receivers, which enables the measurement of both the total and polarised flux of astrophysical objects. These receivers are inherently linearly polarised; however, they can be converted to circular using quarter-wave plates installed on the polarimeter (Marrone and Rao, 2008). These observations came from the SMA Monitoring of AGNs with Polarisation (SMAPOL) programme (Myserlis et al., 2025), which monitors -ray loud blazars, including BL Lac, on a bi-weekly cadence. We calibrated the data using the MIR777https://lweb.cfa.harvard.edu/~cqi/mircook.html package following standard procedures.
@@ 3 VHE intranight variability
In addition to day-to-week variability, we evaluated the VHE -ray intranight variability, a feature less often observed in blazars (other examples are e.g. Aharonian et al., 2007; MAGIC Collaboration et al., 2020). Its detection places significant constrains on the size of the emitting region and on the interpretation of the emission, potentially favouring leptonic models, as hadrons usually require longer acceleration times and have longer cooling times by a factor of . We evaluated the nights with the highest VHE -ray flux: October 20 and November 13. Fig. 2 shows the 5-minute binned light curves.
3.1 October 20 flare
We evaluated the fastest variability timescales observed during the night of October 20. For this night, we considered the integrated flux above 100 GeV, as all data were taken under dark-sky conditions and zenith distances of 20-40∘. The left panel of Fig. 2 shows that the LST-1 data well sample the rise and decay of the flare, varying from 1.0 C.U. to 4.4 C.U. over hours. This well-defined rise and decay profile of the flare allowed us to model the light curve with a rising and decaying exponential function, commonly used for modelling light curves (Norris et al., 1996), including those from blazar flares (see e.g. Aharonian et al., 2007), as
Here, represents the flux of the flare at time , is the baseline flux fixed at cm-2 s-1 from the stable bins at the start and end of the flare, and and are the timescales of the flare rise and decay variations by a factor of . We modelled the 5-minute binned light curve of October 20 as a sum of three possible substructures during the development of the flare, with peak times at min, min, and min, considering the first flux point as min. We also considered the simpler cases of one or two componentsand compared the goodness of fit in each case. We measured values for the triple-, double-, and single-exponential profiles of , , and , which yields significances of and for the triple-exponential model over the double- and single-exponential models, respectively, using Wilks’ theorem (Wilks, 1938). We also verified that adding a fourth component does not improve the fit, showing a preference of only 0.2 over the three-component model. Therefore, we adopted the model based on the sum of three exponential profiles defined by Eq. (1).
We show the fitted function in the left panel of Fig. 2, and we report the calculated rise and decay timescales of the exponential profile in Table 1. These timescales can be used to estimate the doubling flux variability timescale during the rise and decay, which represents the fastest significant variability timescales, as
With this approach, we resolved variability timescales of 18 min with a confidence level of ¿, and detected faster hints between 2.2 min and 4.6 min, albeit with a significance 2.
@@ 3.1 October 20 flare
We also estimated the fastest variability timescales using the doubling-time criterion for flux variations between consecutive temporal bins, following Zhang et al. (1999):
where and are the fluxes at times and . We considered only timescales with variability significance 2. We list the results derived with this approach in the last column of Table 1. This yields a significant doubling flux timescale of min.
Causality arguments imply that very fast variability originates in compact jet regions. The observed fastest variability timescales thus constrain the size of the emitting region responsible for these variations as
where is the radius of the emitting region, represents the fastest variability timescale observed, is the speed of light, is the Doppler factor of the region, and corresponds to the redshift of the source (see Dondi and Ghisellini, 1995; Tavecchio et al., 2010). We adopted a conservative approach, using the shortest timescale with a significance ¿ before trials, namely min. With this timescale and considering a range of Doppler factor values , the size of the emitting region responsible for the fast variability is constrained to cm to cm. The min timescales, if confirmed, imply a region a factor 1.5 to 4 times more compact. Nevertheless, due to their low significance and the typical systematic errors for LST-1 flux measurements (see Abe et al., 2023), these timescales remain unconfirmed.
3.2 November 13 flare
We performed the same analysis for the night of November 13, when we also observed intranight variability in the VHE -ray emission of BL Lac. In this case, owing to the higher energy threshold introduced by moonlight, we considered the flux above 200 GeV. During this night, the emission decayed from 2.8 C.U. to 1.0 C.U. over 1.5 hours. We show the corresponding 5-minute binned light curve in the right panel of Fig. 2. The LST-1 covered only the decay of the VHE -ray flare, which we modelled with a decaying exponential function as
following the form of Eq. (1). We used cm-2 s-1, where the VHE flux appeared to stabilise after the flare. We also report the results for this night in Table 1. Using the estimated timescale min, we again constrained the size of the emitting region with Eq. (4) for typical Doppler factors . This constrains the size between cm and cm. In this case, the doubling flux criterion yields no significant minute-scale variability.
@@ 4 Characterisation of the HE-VHE -ray spectrum
To evaluate possible -ray spectral variability patterns, we studied the evolution of the -ray spectrum during the flare. For this, we applied a Bayesian block (BB) analysis on the VHE -ray light curve using the algorithm developed by Wagner et al. (2022)888https://github.com/swagner-astro/lightcurves, identifying ten BBs of comparable -ray activity with a false-positive rate of 1% (see Fig. 1). Then, using the LAT and LST-1 data, we characterised the high-energy SED peak for the different identified periods from 100 MeV up to a few TeV. We performed a joint fit of the data from both instruments with Gammapy, combining multi-instrument data at the event level, while accounting for instrument response functions (IRFs), backgrounds, and other effects such as EBL absorption (Nievas Rosillo et al., 2025).
@@
For the Fermi-LAT data for each BB, we filtered out all sources in the model with TS¡10. We left free the parameters of all sources within a circular region of 10∘ radius centred at the position of BL Lac and with TS¿30, while for sources within 5∘ of BL Lac with TS¡30, we left only the spectral normalisation as a free parameter. For the remaining sources, we fixed all parameters to the catalogue values. We then performed a pre-fit of the Fermi-LAT data to remove weak sources still included in the model for which the expected counts were ¡5. Finally, we combined the LAT and LST-1 datasets for each BB and calculated the complete -ray spectrum and SED peak, incorporating the response of each instrument into the analysis.
@@
We tested several spectral shapes of the HE-VHE emission, including log-parabola, power-law with exponential cut-off and log-parabola with exponential cut-off spectral models. We find that the first function provides the most accurate fit to our data. Therefore, we used a model consisting of a log-parabola function plus a term accounting for EBL absorption at redshift , following the model from Saldana-Lopez et al. (2021),
@@
To directly compare the different spectra, we adopted GeV for all periods. Table 2 lists the best-fit spectral parameters. Fig. 3 shows an example of the joint SED for the eighth BB (November 13). Figs. 6 and 7 show the remaining SEDs. Due to the minimal variability in spectral curvature, we tested a log-parabola function with fixed at , based on the average value over the entire period. The results were consistent within uncertainties with those reported in Table 2.
@@
The complete characterisation of the high-energy SED peak performed in Sect 4 allowed us to investigate spectral variability patterns during the evolution of the flare. The spectral index at 1 GeV remains stable around , with nearly all BBs consistent within errors at 2. Nevertheless, spectral index variations are particularly noticeable on the nights of October 20 () and November 13 (), indicating that particles are accelerated to considerably higher energies during these flares. This is also reflected in the high-energy peak shift, as reported in Table 2. The IC peak shift is evident for the four brightest BBs, namely the nights of October 20 and 27, and November 13 and 25, which are the only ones with GeV. The average value of is below 2 GeV, while during these nights we observe shifts of up to 5.5 GeV on October 20, 6.5 GeV on November 13, and 2.3 GeV on October 27 and November 25. The Spearman correlation coefficients between and , and between and , indicate a strong correlation ( and , respectively), suggesting a possible harder-when-brighter behaviour. However, this correlation is dominated by the two brightest nights and does not account for uncertainties.
@@ 5 Broadband SEDs
We modelled the evolution of the broadband emission of BL Lac during the high state of 2022. We considered ten broadband SEDs, corresponding to each of the BBs identified above. Fig. 4 shows, as an example, the model for the eighth BB. We present all fitted SED models in Figs. 8 and 9. We note that due to the absence of strictly simultaneous Swift data (0.5 days with respect to the LST-1 observations) to BBs 3, 4, 6, 8, and 9, we adopted a window of 1.5 days for these periods to ensure complete multi-wavelength coverage of the SEDs.
@@
We considered a leptonic-only origin for the broadband emission. We used version v0.4.0 of the python package agnpy (Nigro et al., 2022, 2023). We find that SSC models based on a single emitting region are unable to accurately describe the broadband emission of BL Lac during this period. One-zone models based on EC scattering present an alternative to explain VHE -ray emission and have previously been invoked for BL Lac due to the presence of a small and weak broad line region (BLR), which provides external low-energy photons as seeds for -ray production (MAGIC Collaboration et al., 2019). However, we find that this solution is insufficient to reproduce the broadband emission and variability during the 2022 high state, particularly given the tight constraints on the size of the emitting region imposed by the intranight variability. We therefore considered an SSC model based on two emitting regions (Tavecchio et al., 2011), which has frequently been used to model the multi-wavelength emission of this source (see e.g. MAGIC Collaboration et al., 2019; Sahakyan and Giommi, 2022). We denote the two regions as ‘core’ and ‘blob’, with . We consider the regions to be co-spatial and interacting, motivated by the high correlation of the broadband emission from the optical band up to rays (Appendix D.2). Fig. 9b of MAGIC Collaboration et al. (2019) illustrates a schematic of this layout. Each region is characterised by a magnetic field , a Lorentz factor , a Dopper factor , and an electron population described by a broken power-law distribution with respect to the electron Lorentz factor ,
@@
where is the normalisation of the distribution, and are the minimum and maximum values of the Lorentz factor, and and are the spectral indices below and above the break of the distribution , respectively. We constrained the size of the blob, responsible for the intranight VHE -ray variability, using the estimates presented in Sect. 3. Additionally, we incorporated an EC contribution from external photon fields in the SED model, namely the BLR, the accretion disc, and the dusty torus, since these components have previously been used to reproduce the broad IC SED peak (see e.g. Khatoon et al., 2024). Following Ghisellini and Tavecchio (2009); MAGIC Collaboration et al. (2019), we adopted erg s-1 and cm for the BLR, which implies erg s-1 and cm using the typical relations between BLR, disc, and torus (see Kaspi et al., 2007; Cleary et al., 2007). We assumed a torus covering factor and a BLR reprocessing factor (based on the Hα line) . For the dusty torus, we assumed a typical temperature of 1000 K. We adopted a black hole mass (Ghisellini et al., 2010). We considered the core located at a distance cm from the central engine and the blob at cm. We accounted for possible absorption at GeV-TeV energies and the EBL absorption at VHE. Our analysis shows that absorption due to external radiation fields is negligible.
@@
Multi-zone models of blazar SEDs are characterised by a high degree of degeneracy due to the relatively large number of parameters. This degeneracy generally prevents the use of standard fitting routines based on cost-function minimisation for effectively constraining the parameter space. In such cases, it is necessary to adopt a physically motivated scenario as a guiding framework. In this work, we investigated a scenario in which the evolution of the flare is driven by magnetic reconnection, which is strongly supported by the minute-scale variability detected at VHE with LST-1. To fit the SED of each BB, we used the parameters of the previous block as initial parameters, iterating until we achieved a satisfactory representation of the SED. To reduce the number of free parameters, we kept the magnetic fields , the emitting region sizes , the Lorentz factor , and the Doppler factor fixed, adopting values consistent with the limit values set by the variability analysis. Additionally, we minimised changes in the physical parameters in the core region between adjacent BBs to reflect the slowly evolving nature of the core, unless the time interval between the centre of two adjacent BBs exceeded two days, corresponding to the light-crossing timescale of the core region in our reference frame. In contrast, the blob is considered as a newly formed plasmoid originating from an unresolved magnetic reconnection layer in the core, whose parameters can vary considerably depending on the reconnection layer. Therefore, we based our SED evolution on changes in the electron distribution of each region, namely , , , , , and . The parameter changes were implemented in a manner consistent with the flare evolution timescales. Figs. 8 and 9 show the resulting models for each BB, and Table 3 summarises the corresponding parameters. We discuss the consistency of these models with a magnetic reconnection scenario in Sect. 6.
@@
The leptonic origin of the emission considered here is supported by recent results obtained by the Imaging X-ray Polarimetry Explorer (IXPE) satellite, based on the analysis of the broadband polarised emission of BL Lac (see Middei et al., 2023; Peirson et al., 2023). Within a leptonic framework in which the X-ray spectrum is dominated by the high-energy IC scattering component, the X-ray polarisation degree is expected to be substantially lower than that of the optical band, which originates from electron synchrotron radiation (Liodakis et al., 2019; Peirson and Romani, 2019; Zhang et al., 2024). Middei et al. (2023) measured an upper limit of % for two IXPE observations of BL Lac during May and July 2022, consistent with this scenario given the optical polarisation degree.
@@
Moreover, during strong flares in which the soft X-ray emission has a significant synchrotron radiation contribution, as indicated by a much flatter X-ray spectrum (as in the present case, e.g. BB3 compared to BB8 when the source was at its brightest), we expect a significantly higher polarisation degree of soft X-ray photons than at hard X-ray energies, and even higher than the optical polarisation degree (Liodakis et al., 2019; Peirson and Romani, 2019; Zhang et al., 2024). A third IXPE observation of BL Lac, performed two days after the last LST-1 observation (27-29 November) resulted in a measured X-ray polarisation degree of over the entire X-ray band. After splitting the analysis into two bins corresponding to the soft and hard X-ray regimes, almost all the polarised emission appears in the soft X-ray bin, with a polarisation degree of . Peirson et al. (2023) estimated that the synchrotron emission comprises 44% of the total flux in the soft X-ray band. After accounting only for synchrotron emission and assuming that the remaining 56% of the flux originates from unpolarised SSC emission, these authors calculated the X-ray polarisation fraction to be 49%, much higher than that in optical frequencies (13%), and well within the expectations for an SSC scenario (Peirson and Romani, 2019). Therefore, the leptonic SED modelling performed here is consistent with the expected nature of the broadband emission based on both spectral and multi-wavelength polarisation analyses.